OpenStudy (realitywillslapyou):

What about this? Consider the functions f(x) = 2(x − 7) and g(x) = 2x. What is the value of (f − g)(4)? A. −6 B. 14 C. −14 D. −8

OpenStudy (smokeysthename):

same thing as before plug in 4 for x in both equations then subtract the results. tell me what you get

OpenStudy (realitywillslapyou):

D?

OpenStudy (smokeysthename):

noo okay f(x)=2(x-7) f(4)=2(4-7) f(4)=2(-3) f(4)=-6

OpenStudy (realitywillslapyou):

Oops I typed that wrong letter!

OpenStudy (smokeysthename):

g(4)=2(4) g(4)=8

OpenStudy (smokeysthename):

ohhhh hahaha

OpenStudy (smokeysthename):

well anyways -6-8 = -14 (c) ihope thats the answer you meant to type :P

OpenStudy (realitywillslapyou):

Yeah, there next each other on my keyboard! (˚▽˚)

OpenStudy (smokeysthename):

lol

OpenStudy (realitywillslapyou):

I have a different question up in "Mathematics"

OpenStudy (realitywillslapyou):

Can you try to help me with it?

OpenStudy (smokeysthename):

tag me.

OpenStudy (realitywillslapyou):

Okay!